# Integrable deformations of principal chiral model from

### v1

@@ 1 Introduction
In this paper we consider two-dimensional (or, equivalently 2d or 1+1) integrable field theories of certain type. Let us begin with two widely known examples. The first one is the Landau-Lifshitz model [15, 26]. It is the two-dimensional version of an anisotropic ferromagnetic solid:
@@
where $L(z)$ is the limit of $U(z)$, and $M(z)$ is the limit of $V(z)$. For example, the equation (1.2) in this limit takes the form (with $c_{1}=1$)
@@
In this paper we use construction of integrable Euler-Arnold type tops obtained by a special class of quantum $R$-matrices. In the general case a quantum (non-dynamical) $R$-matrix $R^{\hbar}_{12}(z_{1},z_{2})$ is a solution of the quantum Yang-Baxter equation
@@
Being written in the fundamental representation of ${\rm GL}_{N}$ Lie group $R$-matrix become ${\rm Mat}(N,\mathbb{C})^{\otimes 2}$ valued function of the Planck constant $\hbar$ and the spectral parameters $z_{1},z_{2}$ (in fact, in our case $R^{\hbar}_{12}(z_{1},z_{2})=R^{\hbar}_{12}(z_{1}-z_{2})$). That is
@@
where $\vartheta(z)$ is the first Jacobi theta-function. Summarizing, the quantum $R$-matrix satisfying AYBE (1.11) can be viewed as ${\rm Mat}(N,\mathbb{C})^{\otimes 2}$ valued matrix generalization of the Kronecker elliptic function (1.13), see [22] and [18]. This point of view provides a problem of description of non-commutative generalizations of the elliptic function identities [32].
AYBE can be used for constructing Lax pairs for classical integrable systems of the Euler-Arnold and Gaudin types [17, 19]. For example, one may describe a family of the Euler-Arnold tops (1.7) with the inverse inertia tensor
@@
where ${\rm tr}_{2}$ is a trace over the second tensor component, and $m_{12}(z)$ appears in the quasi-classical expansion of $R^{\hbar}_{12}(z)$ near $\hbar=0$:
@@
In the above consideration we came to some simple finite-dimensional integrable systems by taking the limit $k=0$ in 2d integrable field theories. The inverse problem is to describe integrable field theory generalization for a given finite-dimensional integrable model, where integrability of 1+1 field theory is understood in the sense of the Zakharov-Shabat equation (1.5) and existence of the classical (Maillet type) $r$-matrix structure, while the finite-dimensional integrability is in the Liouville theorem sense through the Lax equation (1.6) and the classical $r$-matrix structure. In the general case this problem is too complicated. At the same time certain families of finite-dimensional integrable systems can be extended to the 1+1 integrable field theories using different approaches. One of them is to use reductions from integrable hierarchies of KP or 2d Toda types, and treat certain time variable in the hierarchy as the space variable $x$. In this way the field analogues of the Calogero-Moser and Ruijsenaars-Schneider models were proposed in [13, 31].
Another approach is based on the affine Higgs bundles by extending the Hitchin approach to finite-dimensional integrable systems [16]. This construction provides 1+1 field analogues of the spin Calogero-Moser model and its multipole generalizations. These type models are non-ultralocal [34], while the models under consideration in this paper are ultralocal. These two type of models are related with each other through 1+1 version of the classical IRF-Vertex correspondence [16, 34, 1]. It is important to mention that the construction based on the affine Higgs bundles explains that $U$-matrix entering the Zakharov-Shabat equation (1.5) has the same form as the Lax matrix $L(z)$ from (1.6). Therefore, we conclude that
@@
There is a third approach to the 2d integrable systems of Costello and Yamazaki, based on the 4d Chern-Simons theory [20]. The works [27, 14] we consider here have been done within this approach. It was demonstrated in [21] that the 4d Chern-Simons construction is essentially coincides with the Affine Higgs bundles description.
@@
The problem is to describe possible integrable deformations of these models. In this paper we study two types of deformations. The first one is given by the so-called twist function. It corresponds to multiplication of $U$-matrix by some function of spectral parameter. We will show that the twist function is naturally identified with the cocentral charge in the affine Higgs bundles. Another set of deformations comes from opportunity to use deformed $r$-matrices in the construction of $U$-matrix similarly to (1.17). In our consideration any $r$-matrix can be used, which comes in the quasi-classical limit (1.15) from a solution of the associative Yang-Baxter equation (1.11) with some special additional properties. The multipole $U$-matrix then takes the form
@@
The paper is organized as follows. First, we recall the general construction underlying Hitchin approach to 2d integrable systems [16]. The Higgs field satisfies the moment map equation, and its solution provides the $U$-matrix of some 2d integrable model. A new result here is a description of possibility to make the cocentral charge to be dependent on the spectral parameter. In the construction of affine Higgs bundles, we identify the twist function with the cocentral charge. Next, we proceed to the associative Yang-Baxter equation. We describe a set of identities underlying Lax equation and the Zakharov-Shabat equation. Then $R$-matrix identities are used to describe 1+1 Gaudin models based on $R$-matrices satisfying AYBE and certain additional properties. Finally, we discuss deformations coming from introducing the twist function. It is closely related to the Gaudin model since it adds new poles to $U$-matrix.
@@ Acknowledgments.
The work of M.A. Olshanetsky was supported by the Russian Science Foundation grant 24-12-00178. The work of Danil Domanevsky was supported by Basis foundation.
2 General construction: Lax operator as affine Higgs connection
2.1 Preliminaries on loop groups and loop algebras.
$\langle X\,,\,Y\rangle=\oint(X,Y)\frac{dy}{y}\,.$
Consider its central extension
$\hat{L}({\mathfrak{g}})=\{(X(x),\kappa)\}\,,~{}~{}\kappa\in{\mathbb{C}}\,.$ (2.1)
The commutator of $\hat{L}({\mathfrak{g}})$ assumes the form
where $[(X_{1},X_{2})]_{0}$ is a commutator on ${\mathfrak{g}}$,
The dual to $\hat{L}({\mathfrak{g}})$ is the Lie coalgebra
$\hat{L}^{*}({\mathfrak{g}})=\{\mathcal{Y}=(Y,{k})\sim({k}\partial+Y)\}\,.$ (2.2)
It is defined through the pairing
$(\mathcal{X},\mathcal{Y})=\langle X,Y\rangle+\kappa{k}\,,$ (2.3)
where ${k}$ is the cocenter.
$L(G)=G\otimes{\mathbb{C}}(y))=\left\{\,\sum_{m}g_{m}y^{m}\,,~{}g_{m}\in G\,\right\}\,.$ (2.4)
$(g,\zeta)\times(g^{\prime},\zeta^{\prime})=\left(gg^{\prime},\zeta\zeta^{\prime}{\mathcal{C}}(g,g^{\prime})\right),$ (2.5)
The adjoint action of $f\in L(G)$ on $\hat{L}({\mathfrak{g}})$ is defined as
${\rm Ad}_{f}\mathcal{X}={\rm Ad}_{f}(X,\kappa)=(fXf^{-1}\,,\,\kappa+\langle f^{-1}\partial f,X\rangle)\,.$ (2.6)
Then the coadjoint action of $L(G)$ assumes the form
${\rm Ad}^{*}_{f}\mathcal{Y}={\rm Ad}^{*}_{f}(Y,{k})=(f^{-1}Yf+{k}f^{-1}\partial f\,,\,{k})\,.$ (2.7)
2.2 Briefly on Hitchin systems
$D_{\bar{A}}=\partial_{\bar{z}}+\bar{A}\,.$ (2.8)
The quotient
$Bun(G)=\{D_{\bar{A}}\}/\mathcal{G}(G)\,,$ (2.9)
where
$\mathcal{G}(G)=C^{\infty}Maps(\Sigma\to G)\,,$ (2.10)
is the moduli space of the holomorphic $G$ bundles over $\Sigma$.
The Higgs field $\Phi$ is a section of the endomorphisms $E\to E\otimes\mathcal{K}_{\Sigma}$, where $\mathcal{K}_{\Sigma}$ is the canonical class. The $G$-Higgs bundle is the pair
$\mathcal{H}=(D_{\bar{A}}\,,~{}\Phi)\,.$ (2.11)
It is the cotangent bundle $T^{*}\mathcal{A}$ to the affine space of holomorphic structures $\mathcal{A}=\{D_{\bar{A}}\}$ on $E$. Therefore the Higgs bundle can be endowed with the canonical symplectic form
$\Omega=\int_{\Sigma}|d^{2}z|(\delta\Phi\stackrel{{\scriptstyle\wedge}}{{,}}\delta\bar{A})\,,$ (2.12)
where (,)(~{},~{})( , ) is a fixed invariant form on the Lie algebra ${\mathfrak{g}}^{\mathbb{C}}$, and $\stackrel{{\scriptstyle\wedge}}{{,}}$ mens the inner product in the symplectic form.
[…]
$D_{\bar{A}}\Phi=0\,.$ (2.13)
This equation defines the holomorphic structure on the Higgs bundle $\mathcal{H}$ (2.11). Let $\mathcal{C}_{0}$ be the set of solutions of (2.13). The symplectic quotient
$\mathcal{M}^{H}(G)=\mathcal{H}//\mathcal{G}(G)\sim\mathcal{C}_{0}/\mathcal{G}(G)$ (2.14)
is the phase space of the Hitchin integrable systems [11, 12, 16] with the Poisson brackets coming from (2.12).
2.3 Affine Hitchin systems
To pass to 2d integrable systems we consider the affine version of the Higgs bundle. For this purpose we replace the group $G$ on the group $\hat{L}(G)$ (2.4), (2.5).
Let us consider an infinite rank $L(G)$-bundle $E^{aff}$ and a line bundle $\mathcal{L}$ over $\Sigma$. Consider also smooth maps $C^{\infty}(\Omega^{(0,1)}(\Sigma)\to\hat{\mathfrak{g}})=\{(\bar{A}(z,\bar{z},y),\bar{\kappa}(z,\bar{z})\}$ (see (2.1)). The components of this map are connection forms on the bundles $E^{aff}$ and $\mathcal{L}$:
$\nabla_{\bar{A}}=\left(\begin{array}[]{c}\partial_{\bar{z}}+\bar{A}(z,\bar{z},y)\\ \partial_{\bar{z}}+\bar{\kappa}(z,\bar{z})\end{array}\right)\,.$ (2.15)
Affine Higgs connection.
Consider the map $C^{\infty}(\mathcal{K}_{\Sigma}\to\hat{L}^{*}({\mathfrak{g}}))$ (2.2) given by
$\Omega=\frac{1}{\pi}\int_{\Sigma}|d^{2}z|\Bigl{(}\langle\delta\Phi\stackrel{{\scriptstyle\wedge}}{{,}}\delta\bar{A}\rangle+\delta{k}\wedge\delta\bar{\kappa}\Bigr{)}\,.$ (2.16)
Marked points.
Introduce two sets of marked points on $\Sigma$
$\mathcal{D}_{1}=\{z_{a}\,,~{}a=1,\ldots,n\}\,,~{}~{}\mathcal{D}_{2}=\{w_{b}\,,~{}b=1,\ldots,m\}\,.$ (2.17)
${\mathcal{O}}(p_{a}^{(0)},c_{a}^{(0)})=\{S^{a}(p_{a}^{(0)},c_{a}^{(0)})=g^{-1}p_{a}^{(0)}g+c_{a}^{(0)}g^{-1}\partial g\,,~{}g\in L(G)\}\,.$ (2.18)
The orbit (passing through a point $p_{a}^{(0)}$, and with the central charge $c_{a}^{(0)}$) is the coset space ${\mathcal{O}}(p^{(0)},c^{(0)})\sim L(G)/H$ for $c^{(0)}\neq 0$, and ${\mathcal{O}}(p^{(0)},0)\sim L(G)/L(H)$, where $H$ is the Cartan subgroup of $G$. Thereby the orbit can be described by the elements of these quotient
$S^{a}\to g_{a}\in L(G)/H\,,~{}{\rm or}~{}g_{a}\in L(G)/L(H)\,.$ (2.19)
The invariants defining the orbit ${\mathcal{O}}(p^{(0)},c^{(0)})$ are the conjugacy classes of the monodromy operator $M$ corresponding to the operator $c^{(0)}\partial+S$. Therefore, there is a one-to-one correspondence between the set of $L(G)$-orbits in the coalgebra $\hat{L}^{*}({\mathfrak{g}})$ (2.2) and the set of conjugacy classes in the group $G$.
Symplectic form.
The symplectic form $\omega_{a}$ on the orbit $\mathcal{O}_{a}$ (2.18) takes the form
$\omega_{a}=\frac{1}{2\pi\imath}\oint\delta(S^{a}(p_{a}^{(0)},c_{a}^{(0)}),g^{-1}\delta g)\frac{dy}{y}\,.$ (2.20)
The corresponding Poisson brackets are
$\{S^{a}_{\alpha}(y),S^{a}_{\beta}(y^{\prime})\}=\delta(y/{y^{\prime}})\sum\limits_{\gamma}c^{\gamma}_{\alpha\beta}S^{a}_{\gamma}(x)+c^{(0)}\kappa_{\alpha\beta}\partial\delta(y/{y^{\prime}})\,,$ (2.21)
where $\kappa_{\alpha\beta}$ is the defined above invariant form on ${\mathfrak{g}}$, $c^{\gamma}_{\alpha\beta}$ are structure constants in ${\mathfrak{g}}$, and $\alpha,\beta,\gamma$ here are indices in some basis in ${\mathfrak{g}}$. The form $\omega$ is invariant under transformations $g\to gf$, $f\in L(G)$. The corresponding moment map is $S^{a}(p_{a}^{(0)},c_{a}^{(0)})$.
Define the symplectic form $\Omega$ on the space of fields
$\mathcal{H}(G)=(\bar{A}\,,\,\Phi\,,\,{k}\,,\,\bar{\kappa}\,,\,\cup_{a=1}^{n}S^{a})\,.$ (2.22)
Taking into account (2.16) and (2.20) we have
$\Omega=\frac{1}{\pi}\int_{\Sigma}|d^{2}z|\Bigl{(}\langle\delta\Phi\stackrel{{\scriptstyle\wedge}}{{,}}\delta\bar{A}\rangle+\delta{k}\wedge\delta\bar{\kappa}-\sum_{a=1}^{n}\omega_{a}\delta(z-z_{a},\bar{z}-\bar{z}_{a})\Bigr{)}\,,$ (2.23)
Gauge symmetries – symplectomorphisms.
Define the map
$G(\Sigma)=C^{\infty}(\Sigma\to L(G))=\{f(z,\bar{z},y)\}\,.$ (2.24)
The gauge group has two components
$\hat{\cal G}:=(G(\Sigma),\mathcal{G}_{1})\,,~{}~{}\mathcal{G}_{1}=\exp\Big{(}\varepsilon_{1}(z,\bar{z})\Big{)}\,,~{}\varepsilon_{1}(z,\bar{z})\in C^{\infty}(\Sigma)$ (2.25)
as well as its Lie algebra
The coalgebras $L^{*}_{0}$ and $L^{*}_{1}$ are distributions on these spaces.
The actions of $L_{0}$ and $L_{1}$ on the fields take the form
$\begin{array}[]{ll}1.&\delta_{\epsilon}\bar{A}(z,\bar{z},x)=-\partial_{\bar{z}}\epsilon+[\epsilon,\bar{A}]\\ 2.&\delta_{\epsilon}\bar{\kappa}(z,\bar{z})=\langle\bar{A}\partial\epsilon\rangle\end{array}$ (2.26)
These formulas are the infinite-dimensional version of (2.6). Also we have
$\begin{array}[]{ll}1.&\delta_{\varepsilon_{1}}\bar{A}(z,\bar{z},x)=0\\ 2.&\delta_{\varepsilon_{1}}\bar{\kappa}(z,\bar{z})=-\partial_{\bar{z}}\varepsilon_{1}\end{array}\,.$ (2.27)
Taking into account the coadjoint action (2.7) we find
$\delta_{\epsilon}\Phi={k}\partial\epsilon+[\Phi,\epsilon]\,,~{}~{}\delta_{\epsilon}{k}=0\,,$ (2.28)
$\delta_{\varepsilon_{1}}\Phi=0\,,~{}~{}\delta_{\varepsilon_{1}}{k}=0\,,$ (2.29)
$\delta_{\epsilon}g_{a}=g_{a}\epsilon(z_{a},\bar{z}_{a})\,,~{}~{}\delta_{\varepsilon_{1}}g_{a}=0\,.$ (2.30)
Symplectic quotient.
The moment maps $\mu_{j}\,:\,\mathcal{H}(G)\to Lie^{*}(\hat{\cal G}^{G})$ are defined by the actions (2.26)-(2.30):
$\mu_{0}=\partial_{\bar{z}}\Phi-{k}\partial\bar{A}+[\bar{A},\Phi]-\sum_{a=1}^{n}S^{a}\delta(z-z_{a},\bar{z}-\bar{z}_{a})\in L_{0}^{*}\,,$ (2.31)
$\mu_{1}=\partial_{\bar{z}}{k}\in L_{1}^{*}=(C^{\infty})^{*}(\Sigma)\,.$ (2.32)
Consider the set of the marked points $\mathcal{D}_{2}$ (2.17). Take
$\mu_{1}=\mu_{1}^{0}\,,~{}~{}\mu_{1}^{0}=\sum_{b=1}^{m}s_{b}\delta(z-w_{b},\bar{z}-\bar{w}_{b})\,,$ (2.33)
where
$\sum_{b=1}^{m}s_{b}=0\,.$ (2.34)
$1.\ \mu_{0}=0\,,\qquad 2.\ \mu_{1}=\mu_{1}^{0}\,.$ (2.35)
In this way we come to the symplectic quotient
This is the set of the gauge equivalent solutions of the moment equations (2.35).

### v4

@@ 1 Introduction
In this paper we consider two-dimensional (or, equivalently 2d or 1+1) integrable field theories of certain type. Let us begin with two widely known examples. The first one is the Landau-Lifshitz model [16, 27]. It is a model of an anisotropic ferromagnetic solid, described by equation
@@
where $L(z)$ is the limit of $U(z)$, and $M(z)$ is the limit of $V(z)$. In particular, the equation (1.2) in this limit takes the form (with $c_{1}=1$)
@@
In this paper we use construction of integrable Euler-Arnold type tops obtained by a special class of quantum $R$-matrices. In the general case a quantum (non-dynamical) $R$-matrix $R^{\hbar}_{12}(z_{1},z_{2})$ is a solution of the quantum Yang-Baxter equation (QYBE)
@@
Being written in the fundamental representation of ${\rm GL}_{N}$ Lie group $R$-matrix become ${\rm Mat}(N,\mathbb{C})^{\otimes 2}$-valued function of the Planck constant $\hbar$ and the spectral parameters $z_{1},z_{2}$ (in fact, in our case $R^{\hbar}_{12}(z_{1},z_{2})=R^{\hbar}_{12}(z_{1}-z_{2})$). That is
@@
where $\vartheta(z)$ is the first Jacobi theta-function. In the general case the QYBE (1.9) and the AYBE (1.11) have different (but intersecting) sets of solutions. At the same time, a certain class of solutions of AYBE satisfying additional properties (see below the skew-symmetry (3.4) and the unitarity (3.5)), satisfies also the QYBE. In this paper we consider exactly these type solutions of AYBE, that is we consider a subset of solutions of QYBE satisfying also AYBE. For this reason we refer to any solution of (1.11) as to $R$-matrix. This is important for our study since AYBE can be used for constructing different type integrable structures (see below) likewise the Fay identity (1.12) is used for constructing the Lax equations and other important objects.
Summarizing, the quantum $R$-matrix satisfying AYBE (1.11) can be viewed as ${\rm Mat}(N,\mathbb{C})^{\otimes 2}$ valued matrix generalization of the Kronecker elliptic function (1.13), see [22] and [19]. This point of view leads to a problem of description of non-commutative generalizations of the elliptic function identities [33].
AYBE can be used for constructing Lax pairs for classical integrable systems of the Euler-Arnold and Gaudin types [18, 20]. In particular, one may describe a family of the Euler-Arnold tops (1.7) with the inverse inertia tensor
@@
where ${\rm tr}_{2}$ is a trace over the second tensor component, and $m_{12}(z)$ is the coefficient appearing in the quasi-classical expansion of $R^{\hbar}_{12}(z)$ near $\hbar=0$:
@@
In the above description we came to some simple finite-dimensional integrable systems by taking the limit $k=0$ in 2d integrable field theories. The inverse problem is to describe integrable field theory generalization for a given finite-dimensional integrable model, where integrability of 1+1 field theory is understood in the sense of the Zakharov-Shabat equation (1.5) and existence of the classical field $r$-matrix structure (of Maillet type), while the finite-dimensional integrability is in the Liouville theorem sense through the Lax equation (1.6) and the classical $r$-matrix structure. In the general case this problem is too complicated. At the same time certain families of finite-dimensional integrable systems can be extended to the 1+1 integrable field theories using different approaches. One of them is to use reductions from integrable hierarchies of KP or 2d Toda types, and treat certain time variable in the hierarchy as the space variable $x$. In this way the field analogues of the Calogero-Moser and Ruijsenaars-Schneider models were proposed in [14, 32].
Another approach is based on the affine Higgs bundles through extending the Hitchin approach to finite-dimensional integrable systems [17]. This construction provides 1+1 field analogues of the spin Calogero-Moser model and its multipole generalizations. These models are non-ultralocal [35], while the models under consideration in this paper are ultralocal. These two type of models are related with each other through 1+1 version of the classical IRF-Vertex correspondence [17, 35, 1]. It is important to mention that the construction based on the affine Higgs bundles explains that $U$-matrix (in the ultralocal case) entering the Zakharov-Shabat equation (1.5) has the same form as the Lax matrix $L(z)$ from (1.6). Therefore, we conclude that
@@
One more approach to the 2d integrable systems is by K. Costello and M. Yamazaki [7]. It is based on the 4d Chern-Simons theory [28, 15]. It was demonstrated in [21] that the 4d construction essentially coincides with the affine Higgs bundles description.
@@
The problem is to describe possible integrable deformations of these models. In this paper we study two types of deformations. The first one is given by the so-called twist function. This deformation corresponds to multiplication of $U$-matrix by some function of spectral parameter. We will show that the twist function is naturally identified with the cocentral charge in the affine Higgs bundles. Another set of deformations comes from opportunity to use deformed $r$-matrices in the construction of $U$-matrix through (1.17). In our description any $r$-matrix can be used, which comes in the quasi-classical limit (1.15) from a solution of the associative Yang-Baxter equation (1.11) with some special additional properties. The multipole $U$-matrix then takes the form
@@
The paper is organized as follows. First, we recall the general construction underlying Hitchin approach to 2d integrable systems [17]. The Higgs field satisfies the moment map equation, and its solution provides the $U$-matrix of some 2d integrable model. A new result here is a description of possibility to make the cocentral charge to be dependent on the spectral parameter. In the construction of the affine Higgs bundles, we identify the twist function with the cocentral charge. Next, we proceed to the associative Yang-Baxter equation. We describe a set of identities underlying Lax equation and the Zakharov-Shabat equation. Then $R$-matrix identities are used to describe 1+1 Gaudin models based on $R$-matrices satisfying AYBE and certain additional properties. Finally, we discuss deformations coming from introducing the twist function. It is closely related to the Gaudin model since it adds new poles to $U$-matrix. In the Appendix we give a very brief review of the Hitchin approach to integrable systems from the group-theoretical viewpoint.
@@ Acknowledgments.
The work of M.A. Olshanetsky was supported by the Russian Science Foundation grant 24-12-00178. The work of Danil Domanevsky was supported by the Foundation for the Advancement of Theoretical Physics and Mathematics ”BASIS”.
2 General construction: Lax operator as connection in the affine Higgs bundle
In this paper we discuss a special (although a wide) class of integrable models described by $U$-matrices, which can be considered as sections of bundles over elliptic curves (or their degenerations) with simple poles (in spectral parameter $z$) at some set of points $z_{1},...,z_{n}$. These models can be described in the framework of the Hitchin approach to integrable systems. Namely, following [17] we define the Lax operator as a section of the affine Higgs bundle which turns into some $U$-matrix of an integrable model after performing symplectic reduction with respect to gauge symmetries. That is, the reduced phase space is the moduli space of the (affine) Higgs bundles. A detailed description of this approach is too long. For this reason we give a brief review of the construction in the Appendix. Main result of the symplectic reduction is that an $U$-matrix satisfies the moment map equation. It can be solved explicitly for some concrete examples of bundles over elliptic curves.
In the finite-dimensional mechanics the Hitchin approach works as follows. The initial phase space is infinite-dimensional. It is a space of pairs $(\Phi,\bar{A})$ including the Higgs field $\Phi$ and the component $\bar{A}$ of the connection defining the holomorphic structure, see (A.11). This space is equipped with the canonical symplectic structure (A.12). By performing the symplectic reduction with respect to the action of gauge group, we arrive to the moment map equation
$\begin{array}[]{c}\displaystyle{\partial_{\bar{z}}\Phi+[\bar{A},\Phi]=\sum_{a=1}^{n}S^{a}\delta(z-z_{a},\bar{z}-\bar{z}_{a})\,.}\end{array}$ (2.1)
A gauge fixation condition makes $\bar{A}$ be either diagonal (an element of the Cartan subalgebra in the Lie algebra of the corresponding Lie group) or equal to zero. Then solutions of (2.1) with respect to $\Phi$ provide Lax matrices of different finite-dimensional integrable systems of classical mechanics. Put it differently, as a result of the reduction $\Phi$ turns into some Lax matrix, see [17, 26].